### v2

# Study on Light Propagation through Space-Time Random Media via Stochastic Partial Differential Equations

Abstract
In this letter, the theory of stochastic partial differential equations is applied to the propagation of light fields in space-time random media. By modeling the fluctuation of refractive index’s square of the media as a random field, we demonstrate that the hyperbolic Anderson model is applicable to describing the propagation of light fields in such media. Additionally, several new quantitative characterizations of the stochastic properties that govern the light fields are derived. Furthermore, the validity of the theoretical framework and corresponding results is experimentally verified by analyzing the statistical properties of the propagated light fields after determining the spatial and temporal stochastic features of the random media. The results presented here provide a more accurate theoretical basis for better understanding random phenomena in emerging domains such as free-space optical communication, detection, and imaging in transparent random media. The study could also have practical guiding significance for experimental system design in these fields.
Introduction.—The problem of light propagation through space-time random media (e.g., the atmosphere and ocean [38, 61, 76, 31, 74]) attracts significant interest in diverse fields, including optical communication and imaging [78, 35, 81, 53, 55, 57], quantum information and sensing [17, 63, 34, 36], random laser [16, 77, 80, 18], adaptive optics [27, 37, 39], photon fluid [66, 22], among others.
The traditional theories of light propagation through random media include the Rytov approximation method [70, 65, 2, 75, 21, 3], the Markov random process approximation method [71, 10], the higher-order moment equation method [14, 9, 42, 32], etc. However, the disparity between the optical wavelength and the characteristic media scale renders existing theories computationally prohibitive for long-distance propagation in realistic random media. Thus, there has developed many simulation methods, exemplary among these simulation methodologies implemented is multiple phase screen model [13, 1] and Monte Carlo method [41, 59].
The aforementioned simulation approaches can partially address specific challenges, such as modeling light intensity distributions under multiple scattering conditions [72, 81, 67] and simulating geometric deformations of beams propagating through turbulence [54, 64, 40]. However, significant limitations persist when dealing with light fields, particularly those propagating through media with strong non-Markovian characteristics [19, 15]. Consequently, developing a comprehensive model capable of characterizing light (especially light fields) transmission in complex space-time random media has become imperative for both theoretical advancements and practical applications.
In this letter, by decomposing the wave equation utilizing an operator separation technique, we derive a stochastic partial differential equation (SPDE) that describes light propagation in space-time random media with enhanced accuracy, which corresponds formally to the hyperbolic Anderson model in SPDE theory [24]. Based on our derivation and recent theoretical advances in the SPDE field [51, 26, 23, 25, 7, 4, 5], we analytically map the media’s stochastic descriptors to the statistical properties of the light field after propagation. Further, outdoor experiments were conducted to verify the validity of these analytical results.
Theory.—The electromagnetic field $u(t,\boldsymbol{x})$ propagating in space-time uniform media satisfies the partial differential equation (PDE) [30]: $\mathcal{L}_{c_{a}}u(t,\boldsymbol{x})=0,$ with $\mathcal{L}_{c_{a}}:=\frac{n^{2}_{a}}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}-\Delta_{\boldsymbol{x}}$, where $\Delta_{\boldsymbol{x}}=\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}$ is the Laplacian operator in dimension three, $n_{a}$ is the refractive index of uniform media, $c$ is the wave speed of light in vacuum. Similar to the potential function applied in time-dependent scattering theory [61, 74], we decompose the square of the refractive index, thereby splitting the operator into two parts:
$\frac{n^{2}(t,\boldsymbol{x})}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}-\Delta_{\boldsymbol{x}}$ $=\frac{\mathbb{E}[n^{2}(t,\boldsymbol{x})]+\mu(t,\boldsymbol{x})}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}-\Delta_{\boldsymbol{x}}$
$=\mathcal{L}_{c_{n}}+\frac{\mu(t,\boldsymbol{x})}{c^{2}}\frac{\partial^{2}}{\partial t^{2}},$ (1)
here $\mu(t,\boldsymbol{x})$ is a space-time random field represents the fluctuation of the square of the media’s refractive index, $\mathcal{L}_{c_{n}}=\frac{\mathbb{E}[n^{2}(t,\boldsymbol{x})]}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}-\Delta_{\boldsymbol{x}}$ is a deterministic two order partial differential operator. Noteworthy, in this letter, $\mathbb{E}[\cdot]$ denotes for ensemble averaging and $\left\langle\cdot\right\rangle$ with respect to average over time. With decomposition Eq.(1), one can rewrite the wave equation as:
$\mathcal{L}_{c_{n}}u(t,\boldsymbol{x})+\frac{\mu(t,\boldsymbol{x})}{c^{2}}\frac{\partial^{2}u(t,\boldsymbol{x})}{\partial t^{2}}=0.$ (2)
Because $\mu(t,\boldsymbol{x})$ is a random field, it is arduous to give solution of Eq.(2) by Sturm–Liouville theory directly [30].
Within the specific context in this letter, there has no strong nonlinear effect in the long-distance propagation process [35, 68], and the time derivative of the phase fluctuations of light fields induced by random media, is significantly lower than the optical frequency. Thus Eq.(2) transforms into the following form:
$\mathcal{L}_{c_{n}}u(t,\boldsymbol{x})$ $={W}(t,\boldsymbol{x})u(t,\boldsymbol{x}),$ (3)
where ${W}(t,\boldsymbol{x}):=\frac{\mu(t,\boldsymbol{x})\omega^{2}}{c^{2}}$, and $\omega$ is the frequency of light field.
The SPDE with the structure given in Eq.(3) is referred to as the hyperbolic Anderson model in the mathematical literature [23]. Accurate modeling of the random field ${W}(t,\boldsymbol{x})$ is essential for solving this model effectively. It is well-established that random media, and notably fully developed turbulence, possess multifractal scaling properties in the spatial domain. [46, 79]. In other words, autocorrelation function of the scalar field in random media has a power-law spectrum as $|\boldsymbol{r}|^{-\gamma}$ [58, 60], here $|\boldsymbol{r}|$ is the Euclidean distance between two points and $\gamma$ is the spatial power law indicator or so-called Extended Self-Similarity (ESS) parameter, the characterization of distinct turbulence is contingent upon the assignment of disparate value to $\gamma$ [69, 60]. In addition, the random media is temporal non-Markovian from the view of time dynamics in most cases [8], the time-domain Markov random media should be treated as a special case. Therefore, we choose fractional Brownian motions (fBm) $B^{H}_{t}$ to describe the correlation of random media in time domain [56, 12]. Based on aforementioned properties, we model the total covariance function of the space-time random media as [62]:
$\mathbb{E}[{W}(t,\boldsymbol{x}),{W}(s,\boldsymbol{y})]=C_{H}(t^{2H}+s^{2H}-|t-s|^{2H})|\boldsymbol{x}-\boldsymbol{y}|^{-\alpha},\quad s,t,\alpha\in\mathbb{R}_{+},\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{3},|\boldsymbol{x}-\boldsymbol{y}|\geq l_{0},$ (4)
where $H\in(0,1)$ is the Hurst index, and $C_{H}$ is a constant related to $H$. The Hurst index $H$ characterizes long-range dependence in stochastic processes:
$\bullet$ $H>1/2$, persistence with positive autocorrelation.
$\bullet$ $H<1/2$, anti-persistence with negative autocorrelation.
$\bullet$ $H=1/2$, memoryless Markov behavior described by standard Brownian motion [12].
And $|\boldsymbol{x}-\boldsymbol{y}|^{-\alpha}$ corresponds to the Riesz kernel function in hyperbolic Anderson model [25]. The parameter $\alpha$ could be determined through experimental measurement, and the restriction $|\boldsymbol{x}-\boldsymbol{y}|\geq l_{0}$ here is to prevent the covariance function from diverging at the zeros, but mathematically it is only necessary for the covariance function to satisfy the Dalang condition [26, 25].
When the covariance function conforms to Eq.(4), there exists a wild solution $u_{w}(t,\boldsymbol{x})$ to the Eq.(2):
$u_{w}(t,\boldsymbol{x})=u_{n}(t,\boldsymbol{x})$ $+\int_{0}^{t}\int_{\mathbb{R}^{3}}G(t-t_{1},\boldsymbol{x}-\boldsymbol{x}_{1})W(dt_{1},d\boldsymbol{x}_{1})+\int_{0}^{t}\int_{\mathbb{R}^{3}}G(t-t_{1},\boldsymbol{x}-\boldsymbol{x}_{1})$
$\times\left(\int_{0}^{t_{1}}\int_{\mathbb{R}^{3}}G(t_{1}-t_{2},\boldsymbol{x}_{1}-\boldsymbol{x}_{2})W(dt_{1},d\boldsymbol{x}_{1})\right)W(dt_{2},d\boldsymbol{x}_{2})+\cdots.$ (5)
Where $t>t_{1}>t_{2}>\cdots>0$, $G(\cdot,\cdot)$ is the Green Function of $\mathcal{L}_{c_{n}}u(t,\boldsymbol{x})=0$, and $u_{n}(t,\boldsymbol{x})$ is the deterministic solution of $\mathcal{L}_{c_{n}}u(t,\boldsymbol{x})=0$ with the same boundary condition and initial value condition [30, 7]. Thus, $u_{w}(t,\boldsymbol{x})$ constitutes a random field in the distributional sense [47]. To further analyze the properties of this solution, we apply Malliavin calculus method, rewriting Eq.(5) as [51, 50]:
$u_{w}(t,\boldsymbol{x})=\sum^{\infty}_{n=0}J_{n}(f_{n}(\cdot,t,\boldsymbol{x})),$ (6)
where $J_{n}$ denotes the $n$-th Wiener-Chaos expansion [7, 50], and $f_{n}$ is the symmetrized distribution in Schwartz Space $\mathcal{S}^{\prime}(\mathbb{R}^{nd})$:
$f_{n}$ $(\cdot,t,\boldsymbol{x})=\frac{1}{n!}\sum_{\pi\in S_{n}}G(t-t_{\pi(n)},\boldsymbol{x}-\boldsymbol{x}_{\pi(n)})$
$\times G(t_{\pi(n)}-t_{\pi(n-1)},\boldsymbol{x}_{\pi(n)}-\boldsymbol{x}_{\pi(n-1)})\cdots$
$\times G(t_{\pi(2)}-t_{\pi(1)},\boldsymbol{x}_{\pi(2)}-\boldsymbol{x}_{\pi(1)})1_{0<t_{\pi(1)}<\cdots<t_{\pi(n)}<t},$ (7)
[…]
For the variance of the intensity $I(t,\boldsymbol{x})=|u_{w}(t,\boldsymbol{x})|^{2}$ of transmitted light field, there has [62]:
$\mathrm{Var}\left[I(t,\boldsymbol{x})\right]=\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j!}\mathbb{E}\left[\left(D^{j}I(t,\boldsymbol{x})\right)^{2}\right],$ (8)
where $D$ is the Malliavin differential operator associates with $W(t,\boldsymbol{x})$, superscript $j$ denotes the Malliavin differential operator acts $j$ times.
We define the strength of the intensity fluctuation as scintillation index $\sigma^{2}$ [11, 53]:
$\sigma^{2}:=\left\langle\Delta I^{2}\right\rangle/\left\langle I\right\rangle^{2}=(\langle I^{2}\rangle/\langle I\rangle^{2})-1.$ (9)
By applying Fubini theorem and the dominated convergence theorem [44, 12], it is straightforward to deduce from Eq.(8) that the scintillation index after propagating a distance $L$ is bounded above by a finite constant [62]:
$\sigma^{2}\big|^{(\infty)}_{L}$ $\simeq\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j!}\int_{S(l_{0})}^{S(L)}j|\boldsymbol{r}|^{-\alpha+2j-2}d\boldsymbol{r}$
$=\int_{S(l_{0})}^{S(L)}\sum_{j=1}^{\infty}\frac{(-1)^{j-1}}{(j-1)!}|\boldsymbol{r}|^{2(j-1)}|\boldsymbol{r}|^{-\alpha}d\boldsymbol{r}$
$=\int_{S(l_{0})}^{S(L)}e^{-|\boldsymbol{r}|^{2}}|\boldsymbol{r}|^{-\alpha}d\boldsymbol{r}$
$\rightarrow\frac{\pi^{3/2}}{2\Gamma(3/2)}\cdot\Gamma\left(\frac{3-\alpha}{2}\right)<\infty,\;\mathrm{as}\;l_{0}\rightarrow 0,L\rightarrow\infty.$ (10)
The upper bound indicates the model we proposed here is able to explain the scintillation saturation phenomenon in experiments [11]. When we take the first $N+1$ $(2\leq N<\infty)$ items in the sequence, the infinite series in formula Eq.(8) is truncated. Further setting $H=1/2$, the discrete version can be obtained by invoking the Lagrange integral mean-value theorem [62]:
$\sigma^{2}\big|^{(N)}_{L}$ $\simeq\sum_{k=1}^{N}\int_{S(l_{k-1})}^{S(l_{k})}\sum_{j=1}^{k}\frac{(-1)^{j-1}}{(j-1)!}|\boldsymbol{r}|^{2(j-1)}|\boldsymbol{r}|^{-\alpha}d\boldsymbol{r}$
$\simeq\sum_{k=1}^{N}V_{k}\sum_{j=1}^{k}\frac{(-1)^{j-1}}{(j-1)!}|\boldsymbol{r}_{k}|^{2(j-1)}|\boldsymbol{r}_{k}|^{-\alpha}.$ (11)
Where $V_{k}:=\int_{S(l_{k-1})}^{S(l_{k})}\mathbf{1}d\boldsymbol{r}$ denotes the volume measure of the integration region over $S(l_{k-1})$ and $S(l_{k})$, and $\{l_{0},l_{1},\cdots,l_{N-1},L,|l_{k-1}<|\boldsymbol{r}_{k}|<l_{k}\}$ is an sequence with elements spaced at the same interval of $(L-l_{0})/N-1$. Furthermore, by regarding each item of Eq.(11) as describing the effect of corresponding phase screen on the scintillation, and take the limit $N\rightarrow\infty$, then we can use the squeeze theorem to explain why the multiple phase screen method could approximately account for the saturation of scintillation when $N\gg 1$ [81, 13, 1].
If aware all the high order moments of the random field $u_{w}(t,\boldsymbol{x})$, we can have a complete description of its overall probability distribution through its generating function. Utilizing the above relationship for lower order moments versus higher order moments, for all integer $p,q\geq 2$ , at the spatial coordinate $\boldsymbol{x}_{0}$, there has [20, 4, 62]:
$\lim_{t\rightarrow\infty}\log\mathbb{E}u_{w}^{p}(t,\boldsymbol{x})$ $=p\left(\frac{p-1}{2}\right)^{\frac{1}{2-\alpha}}C_{1}(\alpha,H)t^{\frac{4H-\alpha}{2-\alpha}},$ (12)
$\dfrac{\langle I^{q}\rangle}{\langle I\rangle^{q}}\bigg|_{\boldsymbol{x}_{0}}$ $=\left(\dfrac{\langle I^{2}\rangle}{\langle I\rangle^{2}}\bigg|_{\boldsymbol{x}_{0}}\right)^{\frac{q(2q-1)^{\frac{1}{2-\alpha}}-q}{2\cdot 3^{\frac{1}{2-\alpha}}-2}},$ (13)
where $C_{1}(\alpha,H)$ is a specific parameter connected with $H,\alpha$ [62]. Formula Eq.(12) is the asymptotic behavior of the solution when the temporal variable $t$ in the hyperbolic Anderson model tends to infinity. As the space power law index $\alpha$ appears explicitly in the expression Eq.(13), the space-time characteristics of the media are reflected in the distribution of the propagated light field.
Subsequently, we analyze the temporal properties of the solution. At a given spatial point, by the isometry property there has [6]:
$\mathbb{E}u_{w}(t,\boldsymbol{x})u_{w}(s,\boldsymbol{x})$ $=C_{H}\int_{\mathbb{R}^{3}}\int_{0}^{t}\int_{0}^{s}dvdw|v-w|^{2H-2}$
$\mathcal{F}G(v-t,\boldsymbol{\xi})\overline{\mathcal{F}G(w-s,\boldsymbol{\xi})}\Lambda(d\boldsymbol{\xi}).$ (14)
here $\Lambda(\cdot)$ denotes the associated Lebesgue measure of $|\boldsymbol{x}|^{-\alpha}$. It is observed that when the Hurst index exceeds 1/2, the covariance function obeys the following identity [6]:
$t^{2H}+s^{2H}-|t-s|^{2H}$
$\equiv 2H(2H-1)\int_{0}^{t}\int_{0}^{s}dvdw|v-w|^{2H-2}.$ (15)
Eq.(14) and Eq.(15) demonstrate that when the Hurst exponent satisfies $H>1/2$, the random field $W(t,\boldsymbol{x})$ exhibits long-range positive temporal dependence. Under this condition, the light field $u_{w}(t,\boldsymbol{x})$ at any fixed spatial location $\boldsymbol{x}_{0}$ likewise displays persistent temporal correlation. Conversely, for $H<1/2$, a negatively correlated random field induces analogous anti-persistent behavior in the resulting light field [58, 12]. This dual dependence is also corroborated by Eq.(12) and Eq.(13), wherein the identity $(4H-\alpha)/(2-\alpha)\equiv 1$ holds exactly at $H=1/2$, marking the transition between correlation regimes.
In practice optical systems, the measured field always represents an averaging of the detected field over the detector area, which can be mathematically described as $F_{R}(t)=\int_{B(R^{2},z_{0})}u_{w}(x,y,z_{0},t)dxdy$, and $B(R^{2},z_{0})$ denotes a disk of radius $R$ at propagation distance $z_{0}$, $R$ is controlled by the entrance pupil of the optical system. By applying the functional central limit theorem [5, 29, 62], we obtain the following upper bound:
$d_{\mathrm{sKL}}\left(Z_{R}(t),\mathcal{N}_{t}(0,1)\right)\leq C_{2}R^{-\frac{\alpha}{2}},$ (16)
here $d_{\mathrm{sKL}}$ denotes the symmetrized Kullback-Leibler (sKL) divergence between two probability density functions [62], $Z_{R}(t):=\{F_{R}(t)-\langle F_{R}(t)\rangle\}/\sqrt{\mathrm{Var}(F_{R}(t))}$, and $\mathcal{N}_{i}(0,1)$ represents a continuous-time sampling from the standard complex Gaussian distributor with zero mean and unit amplitude variance, $C_{2}$ is a constant. Eq.(16) implies that even when $H\neq 1/2$, in the limit of a large spatial integration region, the normalized process $Z_{R}(t)$ converges in distribution to the standard complex Gaussian distribution, with a Hurst index of $1/2$. Moreover, the convergence rate is at least polynomial with exponent $\alpha/2$.
Notably, when $H\neq 1/2$, the light field after propagation exhibits non-Markovian fluctuations under small-aperture detection. Such colored noise is known to be more detrimental to optical communication systems than white noise [78]. The upper bound established in Eq.(16) provides a theoretical foundation for the empirical observation that aperture averaging or spatial diversity reception can effectively mitigate degradation induced by random media in long-distance optical propagation [78, 2, 73, 43, 33].
Figure 1: The schematic diagram depicts the two-point temperature difference measurement device and the large aperture coherent detection system used in this experiment.
Experiment.—To validate the theoretical framework and results proposed in this study, we designed an outdoor experimental setup comprising two functionally integrated systems: a series of impedance-tunable temperature probes to measure the fluctuations of the refractive index square $\mu(t,\boldsymbol{x})$ at two distinct spatial points, and a large-aperture coherent detection system to characterize the amplitude and phase of the propagated light field within the outdoor environment as characterized by measurements of the outdoor random media.
To meet the requirements for experimental validation, key parameters in this design were chosen accordingly. In particular, a laser linewidth of 0.85 kHz was selected to enable precise amplitude and phase measurements over extended distances under realistic conditions, since the phase-stability time of the laser exceeds the fluctuation time of the turbulence-induced random media. An intermediate frequency (IF) of 80 MHz was adopted to ensure sufficient phase resolution within the short-time Fourier transform (STFT) frequency bins. The STFT bin width was set to 10 kHz, which exceeds the Nyquist sampling rate required to capture the characteristic fluctuation scale of the media (approximately 1 kHz [76]), thereby guaranteeing faithful signal reconstruction.
The space-time structure of the atmospheric random media was characterized in situ using an series of TempVue20 Pt100 temperature sensors and a CR1000X datalogger. By systematically varying the inter-sensor spacing, the square of refractive index was derived from the relation $\mu(t,\bm{x})\approx\left[C_{3},p(t,\bm{x})/T(t,\bm{x})+C_{4}\right]^{2}$ [52, 48], where $p$ denotes pressure and $T$ temperature. This enables the computation of the correlation function $K_{\mu}(t_{0},r)=\mathbb{E}[W(t_{0},\bm{0})W(t_{0},\bm{r})]$.
Optically, a laser output is split in a 95:5 ratio into separate signal and local oscillator paths. The signal beam is frequency-shifted by an 80 MHz acousto-optic modulator, expanded to a 300 mm diameter through a Galilean telescope. The beam then propagates over a path length of 588 m in the outdoor atmosphere, over which active tip-tilt correction is applied. At the receiver, a variable-aperture assembly directs the collected light to a balanced detector. The resulting 80 MHz intermediate IF signal is then digitized by a 12-bit ADC and processed in real time using custom algorithms.
Figure 2: (a) Double‑logarithmic plot of the squared refractive index (obtained from outdoor measurements using the TempVue 20 Pt100 sensors) versus separation distance. Experimental data * are compared with two piecewise linear fits (solid and dashed lines). (b1)-(b3) For a 2 mm receiving aperture, the ratios $\frac{\langle I^{q}\rangle}{\langle I\rangle^{q}}\big|_{\boldsymbol{x}_{0}}$ for $q=4,5,6$ are plotted against $\frac{\langle I^{2}\rangle}{\langle I\rangle^{2}}\big|_{\boldsymbol{x}_{0}}$ on double‑logarithmic axes. The results confirm the validity of Eq.(13).
A multi-step process is employed to extract the amplitude and phase of the propagated light fields. Initially, the raw data stream is divided into successive time blocks as: $\{(0,t_{1}),(t_{1},t_{2}),\cdots,(t_{i},t_{i+1}),\cdots\}$. Subsequent Fourier analysis of each block enables the identification of the IF signal’s spectral peak. The complex vector at this peak is then isolated. From this vector, the instantaneous optical field is derived as $A_{i}e^{j\varphi_{i}}$, where the modulus $A_{i}$ and argument $\varphi_{i}$ correspond to the amplitude and phase at the discrete time instance $t_{i}$. It is worth noting that, with such partitioning, experimental measurements using optical fiber delay lines have confirmed that the laser’s intrinsic linewidth is sufficiently narrow, ensuring no significant laser-related additional influences are introduced into the phase measurements. While the intensity and higher-order moments of intensity, as well as the subsequent algorithmic processes, could all be carried out based on the acquired amplitudes and phases.
Following the experiments, we first determined $\mu(t_{0},\boldsymbol{x})$ at a fixed time $t_{0}$ across various propagation distances $r$. The power-law scaling behavior was characterized using a two-segment piecewise fit in double-logarithmic coordinates, minimizing the global residuals. For these measurements, the receiver aperture was set to 2 mm, and the signal after balance detection was sampled at 250 MHz, Fig.2 (b1)-(b3) present the relationship between $\dfrac{\langle I^{q}\rangle}{\langle I\rangle^{q}}\bigg|_{\boldsymbol{x}_{0}}$ and $\dfrac{\langle I^{2}\rangle}{\langle I\rangle^{2}}\bigg|_{\boldsymbol{x}_{0}}$ at $q=4,5,6$, confirming the validity of Eq.(13) as proposed by the model. Comparisons with previous experiments [42, 53] and simulations can also be found in the supplementary materials [62].
Subsequently, the receiver aperture was enlarged from 2 mm to 300 mm in order to examine the validity of Eq.(16), with the data acquisition card operating at its maximum sampling rate of 1 GHz. To quantitatively characterize the scaling behavior, the rescaled range analysis (R/S) method [49, 28, 62] was utilized to estimate the Hurst index associated with fluctuations in both the random media and the phase of light field, as illustrated in Fig.3 (a1)–(a2):
$H_{X}=\lim_{n\rightarrow\infty}\frac{\log{R}/{S}(X_{1},\cdots,X_{n})}{\log n},\;X_{i}=\{\mu(t_{i},\boldsymbol{x}_{0}),\varphi_{i}\}.$ (17)
Figure 3: (a1)-(a2) For a 2 mm aperture, the 24‑hour Hurst index variations of $\mu(t_{i},\boldsymbol{x}_{0})$ and $\varphi(t_{i})$ are strongly correlated (Pearson coefficient 0.905), whereas for a 300 mm aperture the correlation becomes insignificant (Pearson coefficient 0.341). (b) The sKL divergence between $Z_{R}(t)$ and $\mathcal{N}_{t}(0,1)$ is shown on a double‑logarithmic scale as the aperture increases from 2 mm to 300 mm. The dashed line indicates the theoretical upper bound attenuation $C_{2}R^{-\frac{\alpha}{2}}$ in the same coordinates. The results support Eq.(14) and Eq.(16).
Overall, the sKL divergence between $Z_{R}(t)$ and $\mathcal{N}_{t}(0,1)$ was analyzed as a function of increasing aperture diameter, as presented in Fig.3 (b). To conclude, the consistency between experimental measurements and theoretical predictions confirms the validity of the proposed model and its derived results.
Conclusion.—In this letter, we derive a stochastic partial differential equation that models light propagation in space-time random media. The model incorporates temporal fractional Brownian motion and a random field with spatial power-law correlations to capture perturbation processes in non-Markovian transparent random media. Several theoretical properties of the model are analyzed, and outdoor experiments were conducted to validate the proposed approach. Moreover, this work establishes a theoretical foundation for the experimental findings in [53, 42, 62].
The theoretical framework shows direct applicability to critical photonic technologies such as free-space optical communication, coherent imaging through random transparent media, and turbulence-resilient imaging systems. It also holds potential for diagnosing various turbulence regimes via spatial spectral analysis [45], as well as for characterizing types and parameters of random media across practical applications.
Acknowledgments.—This study was carried out under the auspices of National Key Research and Development Program of China (2024YFF0505602).